# Prediction of turbulent channel flow using Fourier neural operator-based machine-learning strategy

### v1

@@ V Conclusions
In the present study, the Fourier-neural-operator (FNO) and implicit U-Net enhanced FNO (IUFNO) are investigated in the large-eddy simulations (LES) of three-dimensional (3D) turbulent channel flows. In the preliminary test, the IUFNO outperforms the FNO in both the training loss and the long-term predictions of velocity statistics. In the a posteriori LES tests, the predictive ability of IUFNO is comprehensively examined and compared against the fDNS benchmark as well as the classical DSM and WALE models at various friction Reynolds numbers, namely $Re_{\tau}\approx 180$, $395$ and $590$.
In comparison to the classical DSM and WALE models, the predicted profiles of mean velocity and the fluctuating velocity magnitude by the IUFNO model are closer to the fDNS results even though the performance at the higher Reynolds number is less satisfying. For the probability density functions (PDFs) of the fluctuating velocity, both the IUFNO and WALE models reconstruct well the fDNS results while the DSM results are much worse.
Additionally, quadrant analysis is performed by calculating the joint PDF of the normalized streamwise and wall-normal fluctuating velocities. The results indicate that the shear event is quite strong in a region close to the wall while it is very weak in the center of the channel, which is also consistent with the shear Reynolds stress profiles. Overall, the IUFNO model gives the best predictions for the joint PDFs of the velocity fluctuations, while its prediction for the Reynolds stress is similar to the WALE model.
Further, the streamwise and spanwise kinetic energy spectra are examined. At all three Reynolds numbers, the IUFNO model gives the best predictions. Finally, the Q-criterion is calculated to examine the vortex structures of the turbulent field. It is observed that the DSM predicts much denser vortex structures compared to the fDNS result. In comparison, the IUFNO and WALE models recover well the vortex structures. These results demonstrate that the IUFNO is a promising framework for the fast prediction of wall-bounded turbulent flows.

### v4

@@ V Conclusions
In the present study, the Fourier-neural-operator (FNO) and implicit U-Net enhanced FNO (IUFNO) are investigated in the large-eddy simulations (LES) of three-dimensional (3D) turbulent channel flows. In the preliminary test, the IUFNO model outperforms the FNO model in both the training loss and the long-term predictions of velocity statistics. In the a posteriori LES tests, the predictive ability of IUFNO is comprehensively examined and compared against the fDNS benchmark as well as the classical DSM and WALE models at various friction Reynolds numbers, namely $Re_{\tau}\approx 180$, $395$ and $590$.
In comparison to the classical DSM and WALE models, the predicted profiles of mean velocity and the fluctuating velocity magnitude by the IUFNO model are closer to the fDNS results even though the performance at the higher Reynolds number is less satisfying. For the probability density functions (PDFs) of the fluctuating velocity, both the IUFNO and WALE models reconstruct well the fDNS results while the DSM results are slightly worse.
Additionally, quadrant analysis is performed by calculating the joint PDF of the normalized streamwise and wall-normal fluctuating velocities. The results indicate that the shear event is quite strong in a region close to the wall while it is very weak in the center of the channel, which is also consistent with the shear Reynolds stress profiles. Overall, the IUFNO model can give quite satisfying predictions for the joint PDFs.
Further, the streamwise and spanwise kinetic energy spectra are examined. At all three Reynolds numbers, the IUFNO model gives the best predictions. Finally, the Q-criterion is calculated to examine the vortex structures of the turbulent field, and the IUFNO model can predict reasonably well the vortex structures. Considering the relatively low computational cost, these results demonstrate that the IUFNO model is a promising framework for the fast prediction of wall-bounded turbulent flows.